OpenStudy (anonymous):

\[\int\limits_{}^{} \arccos(x)dx = \arccos(x)x + \int\limits_{}^{} xdx/\sqrt{1-x^{2}}\] so I used substitution u=x^(2) du/2 = xdx so I have \[\arccos(x) + (1/2)\int\limits_{}^{} du/\sqrt{1-u}\] Then I used trig substitution u = sin^(2)(s) (u)^(1/2) = sin(s) s = arcsin(u^(1/2)) du/ds = 2cos(s)sin(s) du = cos(s)sin(s)ds so I make the substitution \[(2/2)\int\limits_{}^{} \cos(s)\sin(s)ds/\sqrt{1-\sin^{2}} = \int\limits_{}^{} \cos(s)\sin(s)ds/\sqrt{\cos^{2}(s)} = \] \[\int\limits_{}^{} \cos(s)\sin(s)ds/\cos(s) = \int\limits_{}^{} \sin(s)ds = -\cos(s) + c\]

OpenStudy (anonymous):

\[\arccos(x) - \cos(s) + c = \arccos(x) - \cos(\arcsin(x)) + c\] Can you please check my answer?

OpenStudy (experimentx):

the answer is right ... checked from wolfram http://www.wolframalpha.com/input/?i=integrate+arccos+x+dx

OpenStudy (anonymous):

thanks :) I forgot the identities so I didnt k now how to convert it over plus im trying to get the method stuck in my head exams are soon :)

OpenStudy (experimentx):

which plus??

OpenStudy (anonymous):

I think you missinterpreted what I said dont worry about it. Thanks for confirming for me :) I really appreciated it